# Object Detection for Remote Sensing Community

### v1

@@ Introduction
Figure 5: LAE-1M Word Cloud (Part). LAE-COD carries dense categories, while LAE-FOD categories are sparse.
Table 1: LAE-1M dataset contains abundance categories composed of coarse-grained LAE-COD and fine-grained LAE-FOD. LAE-1M samples from these datasets by category and does not count instances of overlap duplicates when slicing.
Datasets Instances Categories
LAE-COD AID (Xia et al. 2017) 34,214 1,380
NWPU-RESISC45 (Hichri 2021) 28,906 1598
SLM (Yuan et al. 2022) 106 1,081
EMS (From Google Earth) 39,013 1,502
LAE-FOD DOTA (Xia et al. 2018) 188,282 18
DIOR (Li et al. 2020) 192,472 20
FAIR1M (Sun et al. 2022b) 1.02 M 5(37)
NWPU VHR-10 (Cheng et al. 2014) 3,651 10
RSOD (Long et al. 2017) 6,950 4
Xview (Lam et al. 2018) $\sim$ 1 M 60
HRSC2016 (Liu et al. 2017) 2,976 1
Condensing-Tower (Zhang and Deng 2019) 2,382 4
@@
Method Backbone Pre-Training Data Fine-Tuning
DIOR($AP_{50}$) DOTAv2.0(mAPmAP)
Generic Object Detection
GASSL (Ayush et al. 2021) ResNet-50 - 67.40 -
CACO (Mall et al. 2023) ResNet-50 Sentinel-2 66.91 -
TOV (Tao et al. 2023) ResNet-50 TOV-NI,TOV-R 70.16 -
Scale-MAE (Reed et al. 2023) ViT-L FMoW 73.81 -
SatLas (Bastani et al. 2023) Swin-B SatlasPretrain 74.10 -
RingMo (Sun et al. 2022a) Swin-B RingMoPretrain 75.90 -
SkySense (Guo et al. 2024) Swin-H multi-modal RSI 78.73 -
MTP (Wang et al. 2024) Swin-H MillionAID 81.10 -
Open-Vocabulary Object Detection
GLIP-FT (Li et al. 2022) Swin-T O365,GoldG,CC3M,SBU 87.8 50.6
GroudingDINO-FT (Liu et al. 2024) Swin-T O365,GoldG,Cap4M 90.4 54.0
GroudingDINO-FT (Liu et al. 2024) Swin-T LAE-1M 91.1 55.1
LAE-DINO-FT (Ours) Swin-T O365,GoldG,Cap4M 92.0 55.5
LAE-DINO-FT (Ours) Swin-T LAE-1M 92.2 57.9
Table 1 summarizes the sub-datasets used for building the LAE-1M dataset. Specifically, for most of the datasets, a 0.4 random sampling rate is adopted if the number of instance of same class us larger than 100. Xview is the only exception, for which we sample 0.2 to eliminate the duplicate instances. The purpose of sampling instances from different classes across all datasets is to maximize the learning of each class’s features while preserving the original dataset’s data distribution. Figure 5 presents a word cloud of a subset of LAE-1M, illustrating that LAE-COD encompasses a richer variety of semantic categories than the LAE-FOD. This diversity aids in improving open-vocabulary modeling for the remote sensing community. Further details on the LAE-Label pipeline can be found in the Appendix Dataset Details section.
Evaluation Benchmarks.
To evaluate the validity of our LAE-1M dataset and LAE-DINO model, DIOR (Li et al. 2020) and DOTAv2.0 (Xia et al. 2018) which are commonly used in the remote sensing community are used as benchmarks as in MTP(Wang et al. 2024). Note that the results of DOTAv2.0 are all based on horizontal detection boxes for building a foundational location detector. In addition, to better validate the open-set detectors, we constructed LAE-80C containing 80 classes as a new remote sensing OVD benchmark. More details are included in Appendix. Based on the above three benchmarks, both the open-set and closed-set detection capabilities are evaluated. Specifically, we introduce the HRRSD (Zhang et al. 2019) dataset with a total of thirteen classes, which contains ten base classes appearing in LAE-1M dataset and three novel classes that do not, to perform the few-shot detection experiments.The $mAP$, $AP_{50}$, and $AP_{75}$ are used as the evaluation metrics.
Implementation Details.
We conducted all pre-training experiments on four A100 GPUs. To avoid memory overflow caused by having too many objects in a single image during batch training, we split image annotations with over 200 objects into smaller groups, ensuring the number of instances remains unchanged. Additionally, the alignment heads’ categories are set to 1600 for open-vocabulary pre-training. During training, key parameters are carefully set: the length of the dynamic vocabulary number in DVC module i.e., $N_{\mathcal{DV}}$ is set to 60, the number of layers $l$ for MDSA and FFN is set to 7, and the hyper-parameters $\alpha$ and $\beta$ of the loss function are set to 1 and 10, respectively. The open-vocabulary pre-training of LAE-DINO lasts approximately 180KK steps, spanning about 48 GPU hours with a batch size of 2 per GPU. More details are provided in More Implementation Details section in Appendix.
Detection Results
Open-Set Detection.
We compare the open-set detection results with two effective OVD methods, GLIP (Li et al. 2022) and GroudingDINO (Liu et al. 2024), trained on natural and remote sensing scenes datasets as shown in Table 2. To train on LAE-1M dataset, we similarly introduce DVC on the GLIP and GroudingDINO. First of all, the detection results find that the OVD method of pre-training on natural scene dataset hardly works on remote sensing open-set detection, indicating a substantial gap between remote sensing and natural scene. Secondly, GroudingDINO has a more powerful open-set detection capability compared to GLIP from DIOR and LAE-80C. Clearly, our LAE-DINO has a better open-set detection compared with GroudingDINO, with an increases of 1.9%, 0.8%, and 2.5% on DIOR, DOTAv2.0, and LAE-80C benchmarks, respectively. These detection results show that our LAE-DINO has a more robust open-set detection in the remote sensing field.
Closed-Set Detection.
To prove the benefits of OVD, we perform fine-tuning experiments in remote sensing scenes, comparing some generic detectors (GD) on DIOR and DOTAv2.0 datasets as shown in Table 4. Most previous GDs are fine-tuned to object detection datasets after pre-training on remote-sensing images using a self-supervised approach. We directly cite the original paper results due to the lack of open source for these generic detectors. For the OVD methods, we also provide results of fine-tuning experiments based on pre-training on natural scene datasets. Comparing the GD and OVD methods, it shows that the OVD method, which introduces textual prompts, is significantly higher than the GD method with a raise of about 6% at least in the $AP_{50}$ on DIOR’s closed-set detection. The results of LAE-DINO fine-tuned on DIOR train set demonstrates that an outstanding performance on the DOTAv2.0 test set, with a $mAP$ of 57.9, an increase of 2.8% compared with GroundingDINO.
Method Fine-Tuning Data DIOR
$AP_{50}$
LAE-DINO DIOR-full 92.2
LAE-DINO DIOR-½ 89.1
LAE-DINO DIOR-¼ 85.6
Ablation Studies
VisGT Analysis.
Table 6: The ablation results on the DIOR test set.
Method Pre-Training Data DIOR
$AP_{50}$
LAE-1M Pre-Training
$PT\text{-}baseline$ LAE-1M 83.6
++ VisGT LAE-1M 85.5
DIOR Fine-Tuning
$FT\text{-}baseline$ - 89.9
++ VisGT - 92.0
++ VisGT LAE-1M 92.2
LAE-1M Analysis.
@@
To explore how LAE-COD and LAE-FOD of LAE-1M work, we set up two sets of comparison experiments on our LAE-DINO as shown in Table 7. We find that the detection of base classes in LAE-FOD can be improved by adding additional LAE-COD for pre-training, where the $mAP$ of DOTAv2.0 test set can be improved by 2.3%. This also implies the feasibility of our LAE-Label to help interpret common categories of remote sensing imagery. As for the annotation quality and novel class detection of the LAE-Label engine, its survey report is in Appendix.
@@
Method Pre-Training Data DIOR DOTAv2.0 LAE-80C
$AP_{50}$ mAPmAP mAPmAP
LAE-DINO LAE-FOD 84.1 44.5 19.1
LAE-DINO LAE-FOD++LAE-COD 85.5(+1.4) 46.8(+2.3) 20.2(+1.1)
VisGT Reanalysis & Visualisation
Conclusion

### v3

@@ Introduction
Datasets Instances
LAE-COD AID (Xia et al. 2017) 34,214
NWPU-RESISC45 (Hichri 2021) 28,906
SLM (Yuan et al. 2022) 106
EMS (From Google Earth) 39,013
LAE-FOD DOTA (Xia et al. 2018) 188,282
DIOR (Li et al. 2020) 192,472
FAIR1M (Sun et al. 2022b) 1.02 M
NWPU VHR-10 (Cheng et al. 2014) 3,651
RSOD (Long et al. 2017) 6,950
Xview (Lam et al. 2018) $\sim$ 1 M
HRSC2016 (Liu et al. 2017) 2,976
Condensing-Tower (Zhang and Deng 2019) 2,382
Table 1: LAE-1M dataset contains abundance categories composed of coarse-grained LAE-COD and fine-grained LAE-FOD.
@@
Method Backbone Pre-Training Data Fine-Tuning
DIOR($AP_{50}$) DOTAv2.0($mAP$)
Generic Object Detection
GASSL (Ayush et al. 2021) ResNet-50 - 67.40 -
CACO (Mall et al. 2023) ResNet-50 Sentinel-2 66.91 -
TOV (Tao et al. 2023) ResNet-50 TOV-NI,TOV-R 70.16 -
Scale-MAE (Reed et al. 2023) ViT-L FMoW 73.81 -
SatLas (Bastani et al. 2023) Swin-B SatlasPretrain 74.10 -
RingMo (Sun et al. 2022a) Swin-B RingMoPretrain 75.90 -
SkySense (Guo et al. 2024) Swin-H multi-modal RSI 78.73 -
MTP (Wang et al. 2024a) Swin-H MillionAID 81.10 -
Open-Vocabulary Object Detection
GLIP-FT (Li et al. 2022) Swin-T O365,GoldG,CC3M,SBU 87.8 50.6
GroudingDINO-FT (Liu et al. 2024b) Swin-T O365,GoldG,Cap4M 90.4 54.0
GroudingDINO-FT (Liu et al. 2024b) Swin-T LAE-1M 91.1 55.1
LAE-DINO-FT (Ours) Swin-T O365,GoldG,Cap4M 92.0 55.5
LAE-DINO-FT (Ours) Swin-T LAE-1M 92.2 57.9
Table 1 summarizes the sub-datasets used for building the LAE-1M dataset. Specifically, for most of the datasets, a 0.4 random sampling rate is adopted if the number of instance of same class us larger than 100. Xview is the only exception, for which we sample 0.2 to eliminate the duplicate instances. The purpose of sampling instances from different classes across all datasets is to maximize the learning of each class’s features while preserving the original dataset’s data distribution.
Evaluation Benchmarks.
To evaluate the validity of our LAE-1M dataset and LAE-DINO model, DIOR (Li et al. 2020) and DOTAv2.0 (Xia et al. 2018) which are commonly used in the remote sensing community are used as benchmarks as in MTP(Wang et al. 2024a). Note that the results of DOTAv2.0 are all based on horizontal detection boxes for building a foundational location detector. In addition, to better validate the open-set detectors, we constructed LAE-80C containing 80 classes as a new remote sensing OVD benchmark. More details are included in Appendix. Based on the above three benchmarks, both the open-set and closed-set detection capabilities are evaluated. Specifically, we introduce the HRRSD (Zhang et al. 2019) dataset with a total of thirteen classes, which contains ten base classes appearing in LAE-1M dataset and three novel classes that do not, to perform the few-shot detection experiments.The $mAP$, $AP_{50}$, and $AP_{75}$ are used as the evaluation metrics.
Implementation Details.
[…]
Detection Results
Open-Set Detection.
We compare the open-set detection results with two effective OVD methods, GLIP (Li et al. 2022) and GroudingDINO (Liu et al. 2024b), trained on natural and remote sensing scenes datasets as shown in Table 2. Note that due to the natural difference of tasks, those CLIP-based (Wang et al. 2024b) and grounding (Mall et al. 2024; Kuckreja et al. 2024; Li et al. 2024a) methods are not considered as competitors. To train on LAE-1M dataset, we similarly introduce DVC on the GLIP and GroudingDINO. First of all, the detection results find that the OVD method of pre-training on natural scene dataset hardly works on remote sensing open-set detection, indicating a substantial gap between remote sensing and natural scene. Secondly, GroudingDINO has a more powerful open-set detection capability compared to GLIP from DIOR and LAE-80C. Clearly, our LAE-DINO has a better open-set detection compared with GroudingDINO, with an increases of 1.9%, 0.8%, and 2.5% on DIOR, DOTAv2.0, and LAE-80C benchmarks, respectively. These detection results show that our LAE-DINO has a more robust open-set detection in the remote sensing field.
Closed-Set Detection.
To prove the benefits of OVD, we perform fine-tuning experiments in remote sensing scenes, comparing some generic detectors (GD) on DIOR and DOTAv2.0 datasets as shown in Table 3. Most previous GDs are fine-tuned to object detection datasets after pre-training on remote-sensing images using a self-supervised approach. We directly cite the original paper results due to the lack of open source for these generic detectors. For the OVD methods, we also provide results of fine-tuning experiments based on pre-training on natural scene datasets. Comparing the GD and OVD methods, it shows that the OVD method, which introduces textual prompts, is significantly higher than the GD method with a raise of about 6% at least in the $AP_{50}$ on DIOR’s closed-set detection. The results of LAE-DINO fine-tuned on DIOR demonstrates that an outstanding performance on the DOTAv2.0, with a $mAP$ of 57.9, an increase of 2.8% compared with GroundingDINO.
Method Fine-Tuning Data DIOR
$AP_{50}$
LAE-DINO DIOR-full 92.2
LAE-DINO DIOR-½ 89.1
LAE-DINO DIOR-¼ 85.6
Ablation Studies
VisGT Analysis.
Method Pre-Training Data DIOR
$AP_{50}$
LAE-1M Pre-Training
$PT\text{-}baseline$ LAE-1M 83.6
$+$ VisGT LAE-1M 85.5
DIOR Fine-Tuning
$FT\text{-}baseline$ 89.9
$+$ VisGT - 92.0
$+$ VisGT LAE-1M 92.2
Table 5: The ablation results on the DIOR test set. PT-baseline denotes the Pre-Training baseline, and FT-baseline denotes the Fine-Tuning baseline.
LAE-1M Analysis.
To explore how LAE-COD and LAE-FOD of LAE-1M work, we set up two sets of comparison experiments on our LAE-DINO as shown in Appendix. We find that the detection of base classes in LAE-FOD can be improved by adding additional LAE-COD for pre-training, where the $mAP$ of DOTAv2.0 test set can be improved by 2.3%. This also implies the feasibility of our LAE-Label to help interpret common categories of remote sensing imagery. As for the annotation quality and novel class detection of the LAE-Label engine, its survey report is in Appendix.
VisGT Reanalysis & Visualisation
Conclusion
Acknowledgments
This work was supported by the National Natural Science Foundation of China (42125503, 42430602). This work also was partially funded by the Ministry of Education and Science of Bulgaria (support for INSAIT, part of the Bulgarian National Roadmap for Research Infrastructure).
Supplementary Material for Locate Anything on Earth: Advancing Open-Vocabulary Object Detection for Remote Sensing Community
More Related Work
Data Engine powered by Large Model.
Large models such as CLIP (Radford et al. 2021) and GPT (Radford et al. 2018) are empowered to change the current paradigm by the Emergent Ability due to the massive amounts of data fed to them for training. In the remote sensing community, multimodal work including a range of CLIP-related work (Liu et al. 2024a; Zhang et al. 2024; Wang et al. 2024b; Pan et al. 2024) has also emerged. Then came a bunch of vision-related large models with data-driven large-scale training, e.g., SAM (Kirillov et al. 2023), InternVL (Chen et al. 2024), etc. Due to the tremendous zero-shot identification ability of the models, some works are based on these models as data engines for automated data labeling. Wang et al. leverage SAM and existing remote sensing object detection datasets to build a pipeline for generating a large-scale remote sensing segmentation dataset SAMRS (Wang et al. 2023). Huang et al. explores the potential application of SAM to process medical images for fine-grained segmentation. As the accuracy requirements of the segmentation task are high, if it is used directly for labeling without adding human checks, it may still be limited in real segmentation.
SAM can sense the exact object edge under the specified point or box prompts but is unaware of the specific category. Large Vision-Language Models (LVLMs), such as CLIP, InternVL, etc., can recognize the relationship between images and text because of their alignment training in large-scale web image-text pair data. Although some work has attempted to give SAM the ability to perceive categories (Chen, Yang, and Zhang 2023), this ability is not mature enough for actual segmentation in remote sensing, and the segmentation quality is degraded. Combined with SAM and LVLMs capabilities, good quality raw labeled data can be obtained.
Prompt-based Object Detection.
Unlike traditional closed-set detection, open-set object detection is a significant change, allowing detector to recognize objects beyond a fixed set of categories with prompt learning. Prompt-based object detection methods can be classified into text-prompted and visual-prompted object detection.
Text-prompted object detection (Gu et al. 2021; Li et al. 2022; Liu et al. 2024b) focuses on guiding visual feature representations by encoding text features that enable them to find the location of regions given textual prompts accurately. This approach often uses a BERT (Devlin et al. 2018) or CLIP (Radford et al. 2021) text encoder as the text backbone. Visual-prompted object detection (Xu et al. 2024; Jiang et al. 2024) mainly defines a visual template as a prompt from the supporting image set to mine the position information of the visual objects. MQ-Det (Xu et al. 2024) incorporates visual prompts from additional supported images with text prompts. However, practical detection is more challenging in obtaining this visual template in advance. We propose the VisGT based on visual-guided text prompt learning, which aims to leverage the visual information further to improve the semantic representation and compensate for the lack of a single text prompt.
Dynamic Vocabulary Strategy.
In natural language processing tasks, the over of the target vocabulary to be processed will not only affect the training speed but also the quality of sentence generation. Wu et al. propose a dynamic vocabulary sequence-tosequence (DVS2S) model which allows each input to possess vocabulary in decoding. Lakew et al. propose a method to transfer knowledge across neural machine translation (NMT) models by means of a shared dynamic vocabulary. Unlike these approaches, our dynamic vocabulary strategy reduces the training vocabulary by randomly selection of samples to address the practical problems in open-vocabulary object detection.
Figure 5: The pipeline of our LAE-Label Engine.
Datasets Image Width Images Instances Categories
$ori.$ $pre.$ $ori.$ $pre.$
LAE-COD AID (Xia et al. 2017) 600 - 10,000 - 34,214 1,380
NWPU-RESISC45 (Hichri 2021) 256 - 31,500 - 28,906 1598
SLM (Yuan et al. 2022) 3,000$\sim$10,001 1,024 22 152 106 1,081
EMS (From Google Earth) 4,864$\sim$11,520 1,024 102 2,605 39,013 1,502
LAE-FOD DOTA (Xia et al. 2018) 800$\sim$4,000 1,024 2,806 17,480 188,282 18
DIOR (Li et al. 2020) 800 - 23,463 - 192,472 20
FAIR1M (Sun et al. 2022b) 1,000$\sim$10,000 600 15,266 64,147 1.02 M 5(37)
NWPU VHR-10 (Cheng et al. 2014) 533$\sim$1,728 - 800 - 3,651 10
RSOD (Long et al. 2017) 512$\sim$1,961 - 976 - 6,950 4
Xview (Lam et al. 2018) 2,576$\sim$5,121 1,024 1,129 26,543 $\sim$ 1 M 60
HRSC2016 (Liu et al. 2017) 300$\sim$1,500 - 1,061 - 2,976 1
Condensing-Tower (Zhang and Deng 2019) 304$\sim$1481 - 892 - 2,382 4
Table 6: LAE-1M Dataset is composed of coarse-grained LAE-COD dataset and fine-grained LAE-FOD dataset. ”-” indicates present ($pre.$) agreement with original ($ori.$). LAE-1M does not count instances of overlap duplicates when slicing.
id area bbox_x0 bbox_y0 bbox_w bbox_h point_input_x point_input_y predicted_iou stability_score crop_box_x0 crop_box_y0 crop_box_w crop_box_h
0 1939 161 210 62 45 197.21875 229.28125 0.9753 0.9574 85 85 171 171
1 593 197 164 35 22 197.21875 175.84375 0.9736 0.9715 85 85 171 171
2 1621 216 208 39 47 229.28125 229.28125 0.9694 0.9639 85 85 171 171
3 153 162 241 14 14 165.15625 250.65625 0.9659 0.9935 85 85 171 171
4 60 185 135 6 8 186.53125 143.78125 0.9605 1.0000 85 85 171 171
5 158 127 248 25 7 143.78125 250.65625 0.9576 0.9568 85 85 171 171
6 44 208 120 5 7 207.90625 122.40625 0.9563 0.9778 85 85 171 171
Table 7: The sample of SAM part for obtaining the RoI of an image.
det_name text class likelihood
cropped_resized_image0_192.0_0.0_589.0_79.0_390.5_39.5.jpg ”Road” with a likelihood of 0.9. The image shows a paved surface with clear lane markings, which is characteristic of a road. The likelihood is not 1 because the image is not clear enough to provide a definitive identification, and there could be other similar-looking features that are not roads. Road 0.9
cropped_resized_image112_0.0_338.0_599.0_479.0_299.5_408.5.jpg ”Airport runway” with a likelihood of 1. The image shows a large, flat, and elongated surface with distinct markings and lines, which are characteristic of an airport runway. The presence of multiple airplanes parked or in motion further supports the identification of this as an airport runway. Airport runway -1
cropped_resized_image113_0.0_0.0_599.0_247.0_299.5_123.5.jpg ”Airport” with a likelihood of 0.9. The image shows a complex network of runways, taxiways, and terminal buildings, which are characteristic features of an airport. The presence of multiple airplanes parked at gates further supports the categorization of this image as an airport. Airport 0.9
cropped_resized_image118_501.0_440.0_599.0_599.0_550.0_519.5.jpg ”Runway” with a likelihood of 0.9. The image shows a large, flat, and elongated area with distinct markings and a pattern that is characteristic of an airport runway. The markings and the way the area is surrounded by other airport infrastructure support this classification. The likelihood is not 1 because the image is not of high resolution, and some details may not be clearly visible. Runway 0.9
cropped_resized_image119_9.0_298.0_247.0_359.0_128.0_328.5.jpg ”Airplanes” with a likelihood of 0.9. The image shows a row of objects that resemble the tail sections of airplanes, which are typically found at airports. The presence of multiple tails in a linear arrangement, along with the context of an airport setting, strongly suggests that these are indeed airplanes. The likelihood is not 1 because the image is not clear enough to confirm the details of the airplanes, such as their make or model, but the overall shape and arrangement strongly indicate that they are airplanes. Airplanes 0.9
cropped_resized_image122_215.0_509.0_356.0_599.0_285.5_554.0.jpg ”Unrecognized” with a likelihood of 0.8. The image is too blurry and lacks clear features to confidently determine the object category. The low resolution and lack of distinct shapes or patterns make it impossible to accurately identify the object. Unrecognized 0.8
Table 8: The sample of LVLM part for annotating the RoI of an image.
More Dataset Details
LAE-Label Engine.
LAE-Label Engine consists of the LAE-FOD dataset construction and the LAE-COD dataset construction. 1) For LAE-FOD dataset construction, we first standardize the dataset annotation format to match the COCO (Lin et al. 2014) format. Then, as there are resolution oversized images in some datasets, e.g., DOTA (Xia et al. 2018), FAIR1M (Sun et al. 2022b) and Xview (Lam et al. 2018), we process to slice the images, as shown in Table 6. In particular, we use the sliced images and annotations from MTP (Wang et al. 2024a), while Xview uses the open-source tool SAHI (Akyon, Altinuc, and Temizel 2022) for slicing to 1024 size. 2) For LAE-COD dataset construction, we first utilize two large-scale datasets for scene classification, AID (Xia et al. 2017) and NWPU- RESIC45 (Hichri 2021) with a low resolution images, to expand remote sensing scenarios. We then take high-resolution imagery from the constituent EMS and SLM (Yuan et al. 2022) datasets from Google Earth [^2] https://earth.google.com/ for cropping. Finally, these raw images are fed to the LAE-COD annotation procedure mentioned in Figure 5. We can also divide it into three parts as follows: a) SAM Part: We first get the region of interest (RoI) by randomly picking points as prompts and then filter the top $K$ objects with the most largest area to be saved as shown in Table 7. b) LVLM Part: We use weights from the InternVL-1.5 version, which are then fed to a specific prompt template “Tell me the possible object category in the remote sensing image by returning a “object category” phrase surrounded by quotation marks and given a likelihood from 0 to 1 “object category” with likelihood, if it is not recognized, output “Unrecognized” and providing reasoning details. ”, which then yields the text for each RoI as shown in Table 8. Empirically, we have found that output categories are more accurate if the LVLM provides the reason for the inference. c) Rlue Part: We first remove some cropped monotone images and categories for “Unrecognized”, and some low likelihood prediction categories. Then, some culling of less accurate categories will be performed.
Figure 6: Raw data labelled by LAE-Label engine without rule-based filtering. The preliminary labelling results have already some accuracy and can provide a boost to the LAE model to understand Earth.
Figure 6 illustrates the raw image without rule-based filter annotation. The rule-based filtering approach employs the removal of monotonous images and the culling of some irrelevant and homogeneous categories.
Datasets Original Category Number Original Categories Selected Category Number Selected Categories
DOTA 18 plane, ship, storage tank, baseball diamond, tennis court, basketball court, ground track field, harbor, bridge, large vehicle, small vehicle, helicopter, roundabout, soccer ball field, swimming pool, container crane, airport, helipad 6 helicopter, roundabout, soccer ball field, swimming pool, container crane, helipad
DIOR 20 airplane, airport, ground track field, harbor, baseball field, overpass, basketball court, ship, bridge, stadium, storage tank, tennis court, expressway service area, train station, expressway toll station, vehicle, golf course, wind mill, chimney, dam 18 airplane, airport, ground track field, harbor, baseball field, overpass, basketball court, bridge, stadium, storage tank, tennis court, expressway service area, train station, expressway toll station, vehicle, golf course, wind mill, dam
FAIR1M 5(37) airplane (Boeing 737, Boeing 777, Boeing 747, Boeing 787, Airbus A320, Airbus A220, Airbus A330, Airbus A350, COMAC C919, and COMAC ARJ21), ship (passenger ship, motorboat, fishing boat, tugboat, engineering ship, liquid cargo ship, dry cargo ship, warship), vehicle (small car, bus, cargo truck, dump truck, van, trailer, tractor, truck tractor, excavator), court(basketball court, tennis court, football field, baseball field), road (intersection, roundabout, bridge) 18 passenger ship, motorboat, fishing boat, tugboat, engineering ship, liquid cargo ship, dry cargo ship, warship, small car, bus, cargo truck, dump truck, van, trailer, tractor, truck tractor, excavator, intersection
Xview 60 Fixed-wing Aircraft, Small Aircraft, Cargo Plane, Helicopter, Passenger Vehicle, Small Car, Bus, Pickup Truck, Utility Truck, Truck, Cargo Truck, Truck w/Box, Truck Tractor, Trailer, Truck w/Flatbed, Truck w/Liquid, Crane Truck, Railway Vehicle, Passenger Car, Cargo Car, Flat Car, Tank car, Locomotive, Maritime Vessel, Motorboat, Sailboat, Tugboat, Barge, Fishing Vessel, Ferry, Yacht, Container Ship, Oil Tanker, Engineering Vehicle, Tower crane, Container Crane, Reach Stacker, Straddle Carrier, Mobile Crane, Dump Truck, Haul Truck, Scraper/Tractor, Front loader/Bulldozer, Excavator, Cement Mixer, Ground Grader, Hut/Tent, Shed, Building, Aircraft Hangar, Damaged Building, Facility, Construction Site, Vehicle Lot, Helipad, Storage Tank, Shipping container lot, Shipping Container, Pylon, Tower 34 Fixed-wing Aircraft, Small Aircraft, Cargo Plane, Pickup Truck, Utility Truck, Passenger Car, Cargo Car, Flat Car, Locomotive, Sailboat, Barge, Ferry, Yacht, Oil Tanker, Engineering Vehicle, Tower crane, Reach Stacker, Straddle Carrier, Mobile Crane, Haul Truck, Front loader/Bulldozer, Cement Mixer, Ground Grader, Hut/Tent, Shed, Building, Aircraft Hangar, Damaged Building, Facility, Construction Site, Shipping container lot Shipping Container, Pylon, Tower
Condensing-Tower 4 working chimney, unworking chimney, working condensing tower, unworking condensing tower 4 working chimney, unworking chimney, working condensing tower, unworking condensing tower
Table 9: LAE-80C is sampled from the validation set of multiple remote sensing object detection datasets to filter the categories that are as semantically non-overlapping as possible.
LAE-1M Word Cloud.
Figure 7 presents a word cloud of a subset of LAE-1M, illustrating that LAE-COD encompasses a richer variety of semantic categories than the LAE-FOD. This diversity aids in improving open-vocabulary modeling for the remote sensing community.
Figure 7: LAE-1M Word Cloud (Part). LAE-COD carries dense categories, while LAE-FOD categories are sparse. Although dense categories include unbounded ones like roads, vegetation, and water, sparse categories are better understood by learning this part.
LAE-80C Benchmark.
In order to expand the number of classes in the DIOR (Li et al. 2020) and DOTAv2.0 (Xia et al. 2018) datasets, we have introduced additional classes selected from various sub-datasets’ test sets. This was done to minimize class duplication, as depicted in Table 9. We prioritized preserving the classes from the high-quality DIOR dataset and then added the categories not available in DIOR from the DOTAv2.0 dataset. For more detailed category data, we included categories from FAIR1M (Sun et al. 2022b) and Xview (Lam et al. 2018) datasets. Additionally, we incorporated the categories of different states in the Condensing-Tower (Zhang and Deng 2019) dataset. In the end, we combined these categories to create a benchmark with 80 categories.
More Implementation Details
LAE-Label Engine.
For the SAM model, we use a version of SAM-ViT-H, along with a series of hyperparameters in the execution, where we take 32 points randomly for each image. Then, the threshold for IOU is set to 0.86, the threshold for score is set to 0.92, and the downsampling factor is set to 2. For images cropped to 1024 resolution size, we take the top ten RoIs with the largest area, while for small resolution images, the top five RoIs with the largest area are used. For the InternVL model, we use the InternVL-1.5 version, which performs parallel inference on eight A100s and completes inference on a single image in about 1.5 on average.
LAE-DINO Model.
[…]
More Results
Few-shot Detection.
@@ Acknowledgments
LAE-COD Quality Evaluation.
To further verify the LAE-Label engine’s annotation quality, we randomly sampled ten image annotations from LAE-COD. Then, we provided some users a quality score (1-5 points) from category accuracy and box accuracy, respectively. The results of the survey are shown in Figure 8. This survey shows that the overall quality of annotation is good, and the quality of classification is generally better than the quality of border. Most adopt a positive attitude towards the accuracy and completeness of the labeling but still need to pay attention to the further improvement of the border quality.
Figure 8: The result of a randomly ten selected sample is scored out of 5 for the box accuracy and the category accuracy and out of ten for the overall accuracy.
LAE-1M Reanalysis.
@@
To explore how LAE-COD and LAE-FOD of LAE-1M work, we set up two sets of comparison experiments on our LAE-DINO as shown in Table 11. We find that the detection of base classes in LAE-FOD can be improved by adding additional LAE-COD for pre-training, where the $mAP$ of DOTAv2.0 test set can be improved by 2.3%. This also implies the feasibility of our LAE-Label to help interpret common categories of remote sensing imagery.
Method Pre-Training Data DIOR DOTAv2.0 LAE-80C
$AP_{50}$ $mAP$ $mAP$
LAE-DINO LAE-FOD 84.1 44.5 19.1
LAE-DINO LAE-FOD$+$LAE-COD 85.5(+1.4) 46.8(+2.3) 20.2(+1.1)
@@
The LAE-Label engine identifies common categories in remote sensing object detection and enhances the detection of some novel categories. Figure 9 shows the detection results of the LAE-COD dataset on the DIOR common classes and some uncommon novel classes. Figure 9(a) represents the zero-shot detection results on the DIOR test set only after LAE-COD training, where $AP_{50}$ of the category is the top 10 category. Figure 9(b) shows that the detection results of some rare classes are pre-trained with or without LAE-COD, which cannot be quantitatively evaluated because there is no corresponding data set, and the data was obtained from Google Earth. We can find that this semi-automated labeling approach LAE-Label significantly improves the shortcomings of existing remote sensing object detection datasets. Additionally, including more categories offers a feasible solution for the LAE task.
Figure 9: The role of LAE-COD dataset in common class detection and novel class detection. The source of the image to be detected is from Google Earth.
VisGT Reanalysis.
Inspired by some remote sensing image-text retrieval work(Pan et al. 2023b), we visualise the distribution of image features for VisGT. Figure 10 shows the detection results on the DIOR test set with different $\beta$ values of $\mathcal{L}_{VisGT}$. We also visualized the visual interpolation of VisGT at different $\beta$, using the t-SNE (Van der Maaten and Hinton 2008) method. The clustered red squares represent the visual features of the image as a whole in semantic space. It can be observed that imposing constraints on VisGT will be better than the prompts without constraints, but over-imposing constraints can also affect the detection results. VisGT works best, especially when $\beta$ is around 10. VisGT can control $\beta$ of $\mathcal{L}_{VisGT}$ to achieve engagement with visual-guided textual prompts, supplementing insufficient textual information to influence the final detection results. In complex remote sensing scenes, VisGT can first retrieve the approximate scene by mapping the visual features into the semantic space and then perform the detection.
Figure 10: The detection results on the DIOR test set with different $\beta$ values of $\mathcal{L}_{VisGT}$, which is visualised the visual interpolation of VisGT at different $\beta$, using the t-SNE (Van der Maaten and Hinton 2008) method.
Visualization.
We visualize open-set detection for GLIP (Li et al. 2022), GroundingDINO (Liu et al. 2024b), and LAE-DINO, as shown in Figure 11. We select a few common categories, such as tennis court, ship and harbor, and expressway service area. Through observation, it is found that GLIP results are biased towards finding more that may be similar, while GroundingDINO finds it as accurate as possible, even if it misses some. Our LAE-DINO detection results are superior to the other two methods. It also side-steps the effectiveness of retrieving scenarios before the detection approach.
Figure 11: Visualization of GLIP, GroudingDINO and LAE-DINO in open-set detection, both pre-trained on LAE-1M dataset.